# DeepQuark: deep-neural-network approach to multiquark bound states

### v1

@@ IV Results and Discussions
$ccqq\bar{c}$ $\frac{1}{2}^{-},\frac{3}{2}^{-}$ $\eta_{c}\Lambda_{c},J/\psi\Lambda_{c}$ NB $\sim$350% $\sim$65%
$\frac{5}{2}^{-}$ $\bar{D}^{*}\Xi_{cc}^{*}$3 27% 73% 0% 0.50 1.39 1.90 1.73 1.38
$bbqq\bar{b}$ $\frac{1}{2}^{-},\frac{3}{2}^{-}$ $\eta_{b}\Lambda_{b},\Upsilon\Lambda_{b}$ NB $\sim$350% $\sim$65%
$\frac{5}{2}^{-}$ $B^{*}\Xi_{bb}^{*}$14 19% 80% 1% 0.30 0.89 1.22 0.88 0.88
@@ V Conclusions and Outlooks
For the first time, we implement the DNN-based VMC approach for the multiquark bound states, whose complexity surpasses that of electron or nucleon systems due to strong SU(3) color interactions. We design a novel and high-efficiency architecture, DeepQuark, to address the unique challenges in multiquark systems like stronger correlations, extra discrete quantum numbers, and intractable confinement interaction. We have shown that DeepQuark is competitive with state-of-the-art approaches including DMC and GEM in baryon and tetraquark systems, reaching high accuracies on ground-state energies with low computational costs. In the nucleon, DeepQuark easily adapts both two-body and flux-tube interactions. In tetraquark systems, it consistently describe meson molecule $T_{cc}$ and compact tetraquark $T_{bb}$ with an unbiased form of wave function ansatz. Moreover, we use DeepQuark to investigate triply heavy pentaquark ($QQqq\bar{Q}$) systems. We obtain weakly bound $\bar{D}^{*}\Xi_{cc}^{*}$ molecule $P_{cc\bar{c}}(5715)$ with $S=\frac{5}{2}$ and its bottom partner $P_{bb\bar{b}}(15569)$. They can be viewed as the analogs of the molecular $T_{cc}$. We recommend experimental search of $P_{cc\bar{c}}(5715)$ in the D-wave $J/\psi\Lambda_{c}$ channel.
@@
DeepQuark holds great promise for extension to diverse pentaquark and even hexaquark systems, overcoming the computational barriers that limit conventional methods. Such investigations can provide forward-looking predictions for future experiments. Furthermore, DeepQuark serves as a powerful framework for exploring confining mechanism beyond two-body interactions in tetraquark states, which may offer valuable insights into multiquark inner structures and nonperturbative QCD. The techniques employed in DeepQuark could enrich the toolkit of deep learning approaches integrated with physics, particularly shedding new light on quantum many-body phenomena.
Acknowledgment
@@
We thank Yao Ma, Yan-Ke Chen and Liang-Zhen Wen for the helpful discussions. L.M. also thank Yilong Yang and Pengwei Zhao for discussions on DNN and VMC. This project was supported by the National Natural Science Foundation of China (No. 12475137), and ERC NuclearTheory (Grant No. 885150). The codes are developed based on the NetKet package [72]. The computational resources are supported by High-performance Computing Platform of Peking University.
DATA AVAILABILITY
@@
The data supporting this study’s findings are available within the article.
Appendix A Quark Model
The parameters of the AL1 potential [47, 48] adopted in this work are taken from Ref. [48] and listed in Table 3. The masses of heavy mesons and baryons in the model are given in Table 4.
Table 3: The parameters in the AL1 quark potential model.
@@ Appendix A Quark Model
0.5069 0.1653 0.8321 1.8609 5.227 1.836 0.315 $A\left(\frac{2m_{i}m_{j}}{m_{i}+m_{j}}\right)^{-B}$ 1.6553 0.2204 2.02×$\times 10^{-3}$
Table 4: The masses (in MeV) of heavy mesons and baryons in the AL1 quark potential model. The results from Gaussian expansion method and the experimental values are shown for comparison.
@@
For the nucleon, we also test the flux-tube confinement model as shown in Fig. 3, where the linear pairwise confinement interactions $V_{\text{conf},ij}$ in Eq. (2) are replaced by the flux-tube interaction in Eq. (3). The string tension $\sigma$ in Eq. (3) is given as $\sigma=0.9204\lambda$ [58], which was determined by fitting the experimental mass of $\Omega^{-}(sss)$ and is consistent with the best fitting parameters in lattice QCD simulation [52]. In baryon systems, the minimal length of color flux tubes can be solved analytically [53] and is given as
$L_{\text{min}}=\begin{cases}\left[\frac{1}{2}(a^{2}+b^{2}+c^{2})+\frac{\sqrt{3}}{2}\sqrt{d(d-2a)(d-2b)(d-2c)}\right]^{1/2}&\max(\theta_{a},\theta_{b},\theta_{c})<\frac{2\pi}{3},\\ d-\max(a,b,c)&\max(\theta_{a},\theta_{b},\theta_{c})>\frac{2\pi}{3},\end{cases}$ (11)
@@
Figure 3: Two confinement scenarios for the baryons. The left and right panels represent the pairwise confinement mechanism and the flux-tube confinement mechanism, respectively.
Appendix B Details of the optimization process

### v2

@@ Abstract
$ccqq\bar{c}$ $\frac{1}{2}^{-},\frac{3}{2}^{-}$ $\eta_{c}\Lambda_{c},J/\psi\Lambda_{c}$ NB $\sim$35% 0% $\sim$65%
$\frac{5}{2}^{-}$ $\bar{D}^{*}\Xi_{cc}^{*}$ -3 27% 73% 0% 0.50 1.39 1.90 1.73 1.38
$bbqq\bar{b}$ $\frac{1}{2}^{-},\frac{3}{2}^{-}$ $\eta_{b}\Lambda_{b},\Upsilon\Lambda_{b}$ NB $\sim$35% 0% $\sim$65%
$\frac{5}{2}^{-}$ $B^{*}\Xi_{bb}^{*}$ -14 19% 80% 1% 0.30 0.89 1.22 0.88 0.88
@@
Conclusions and outlooks—We introduce a novel and high-efficiency DNN-based VMC approach, DeepQuark, to solving the multiquark bound states. We exploit the expressive power of DNN to construct an unbiased and symmetry-informed wave function ansatz, which is capable of handling complex interactions and describing various multiquark configurations. DeepQuark is competitive with state-of-the-art approaches including DMC and GEM in baryon and tetraquark systems, reaching high accuracies on ground-state energies with low computational costs. In the nucleon, DeepQuark easily adapts both two-body and flux-tube interactions. In tetraquark systems, it consistently describes meson molecule $T_{cc}$ and compact tetraquark $T_{bb}$. Moreover, we employ DeepQuark to investigate triply heavy pentaquark ($QQqq\bar{Q}$) systems. We obtain weakly bound $\bar{D}^{*}\Xi_{cc}^{*}$ molecule $P_{cc\bar{c}}(5715)$ with $S=\frac{5}{2}$ and its bottom partner $P_{bb\bar{b}}(15569)$. They can be viewed as the analogs of the molecular $T_{cc}$. We recommend experimental search for $P_{cc\bar{c}}(5715)$ in the D-wave $J/\psi\Lambda_{c}$ channel, which could be explored in future LHCb experiments.
@@
DeepQuark holds great promise for extension to diverse pentaquark and even hexaquark systems, overcoming the computational barriers that limit conventional methods. Such investigations can provide forward-looking predictions for future experiments. Furthermore, DeepQuark is capable of incorporating a variety of potential models with different dynamical ingredients, including complex interactions beyond two-body forces. It therefore provides a powerful framework for investigating lattice-inspired potentials and exploring the confining mechanisms of tetraquark states in future studies, potentially offering valuable insights into the multiquark internal structures and nonperturbative QCD. The techniques employed in DeepQuark could enrich the toolkit of deep learning approaches integrated with physics, particularly shedding new light on quantum many-body phenomena.
@@
Acknowledgments—We thank Yao Ma, Yan-Ke Chen, and Liang-Zhen Wen for the helpful discussions. L.M. also thanks Yilong Yang and Pengwei Zhao for discussions on DNN and VMC. This project was supported by the National Natural Science Foundation of China (Grant No. 12475137 and No.125B2105), ERC NuclearTheory (Grant No. 885150) and Start-up Funds of Southeast University (Grant No. 4007022506). The codes are developed on top of JAX [73], Flax [74], and the NetKet package [75, 76]. The computational resources are supported by High-performance Computing Platform of Peking University.
@@
Data Availability—The data supporting this study’s findings are available within the article. The codes that can reproduce the reported results are available on GitHub [77].
Supplemental Materials: DeepQuark: deep-neural-network approach to multiquark bound states
This supplemental material provides additional details on the quark potential model, neural-network optimization process and results of electron systems.
Supplemental Material A Quark Potential Models
A.1 The AL1 potential and flux-tube potential
In a nonrelativisitic quark potential model, the Hamiltonian of a multiquark system in the center-of-mass frame reads,
$H=\sum_{i=1}^{n}\left(m_{i}+\frac{p_{i}^{2}}{2m_{i}}\right)+\sum_{i<j=1}^{n}V_{ij}.$ (S1)
In this work, we adopt the AL1 potential [50, 51], consisting of the one-gluon-exchange (OGE) interaction and two-body linear confinement interaction,
$V_{\text{OGE},ij}=-\frac{3}{16}\boldsymbol{\lambda}_{i}\cdot\boldsymbol{\lambda}_{j}\left(-\frac{\kappa}{r_{ij}}-\Lambda+\frac{8\pi\kappa^{\prime}}{3m_{i}m_{j}}\frac{e^{-r_{ij}^{2}/r_{0}^{2}}}{\pi^{3/2}r_{0}^{3}}\boldsymbol{s}_{i}\cdot\boldsymbol{s}_{j}\right),$ (S2)
For baryon systems, an additional phenomenological term $V_{123}=-\frac{C}{m_{1}m_{2}m_{3}}$ is introduced to mimic the three-body interaction effect [51]. For hadron systems with heavy quarks, this term is suppressed by the heavy quark mass and therefore is neglected in the calculations. The parameters of the AL1 potential are taken from Ref. [51] and listed in Table SM-I. The masses of heavy mesons and baryons in the model are given in Table SM-II.
Table SM-I: The parameters in the AL1 quark potential model.
@@ A.1 The AL1 potential and flux-tube potential
[…]
Table SM-II: The masses (in MeV) of heavy mesons and baryons in the AL1 quark potential model. The results from Gaussian expansion method and the experimental values are shown for comparison.
@@
Figure S1: Two confinement scenarios for the baryons. The left and right panels represent the pairwise confinement mechanism and the flux-tube confinement mechanism, respectively.
For the nucleon system, we also test the flux-tube confinement potential as shown in Fig. S1, where the linear pairwise confinement interactions $V_{\text{conf},ij}$ are replaced by the flux-tube interaction $V^{\text{ft}}_{\text{conf}}=\sigma L_{\text{min}}$. The string tension $\sigma$ is given as $\sigma=0.9204\lambda$ [59], which was determined by fitting the experimental mass of $\Omega^{-}(sss)$ and is consistent with the best fitting parameters in lattice QCD simulation [55]. In baryon systems, the minimal length of color flux tubes can be solved analytically [56] and is given as
$L_{\text{min}}=\begin{cases}\left[\frac{1}{2}(a^{2}+b^{2}+c^{2})+\frac{\sqrt{3}}{2}\sqrt{d(d-2a)(d-2b)(d-2c)}\right]^{1/2}&\max(\theta_{a},\theta_{b},\theta_{c})<\frac{2\pi}{3},\\ d-\max(a,b,c)&\max(\theta_{a},\theta_{b},\theta_{c})>\frac{2\pi}{3},\end{cases}$ (S3)
@@
A.2 Uncertainty of the AL1 model
Generally, the uncertainty of a phenomenological model consists of two parts: the discrepancy between the model and the underlying theory and the uncertainty of the model parameters.
The discrepancy between the model and the underlying theory: The AL1 model include some core ingredients of QCD, such as the short-range one-gluon-exchange interaction and the long-range confinement. However, given the non-perturbative nature of QCD in low-energy regime, there are still many unknown factors in the quark-quark interactions, such as the effects of loop corrections. These effects may be partly compensated by fitting the model parameters to reproduce the experimental meson spectra. However, some factors may be missing in the model, and it is highly challenging to quantify such uncertainties, as it would require full knowledge of the non-perturbative QCD. We can only infer from the meson spectra that the energy results of the quark potential model may deviate from the experimental values by around tens of MeV.
The uncertainty of the model parameters: The parameters in the AL1 model include four constituent quark masses and six potential parameters. They were determined by fitting the meson spectra across all flavor sectors [51]. Given the high dimensionality of the parameter space and the strong correlations among parameters, a comprehensive determination of parameter uncertainties would require an extensive exploration of the full ten-dimensional parameter space, which is highly computationally demanding. Instead, we employ a sampling-and-selection strategy as a practical approach to provide a reasonable estimate of parameter uncertainties. We vary the parameters within a range around their best-fit values, such that the goodness of fit of the meson spectra remains comparable.
To be more specific, starting from the original parameter set, we generate new parameter sets by sampling each parameter independently from a Gaussian distribution centered at its original value, with a standard deviation equal to 0.5% of that value. For each generated parameter set, we compute the corresponding mass spectra of sixteen ground-state mesons, including $\pi,\rho,K,K^{*},D,D^{*},D_{s},D_{s}^{*},B,B^{*},B_{s},B_{s}^{*},\eta_{c},J/\psi,\eta_{b},\Upsilon$, and calculate the $\chi^{2}$ with respect to the experimental data,
$\chi^{2}=\sum_{i=1}^{16}\left(\frac{m_{i}^{\text{Theo.}}-m_{i}^{\text{Exp.}}}{\Delta m_{i}}\right)^{2},$ (S4)
where $\Delta m_{i}$ is taken to be $0.01m_{i}^{\text{Exp.}}$ to account for the model uncertainty. We then compare the $\chi^{2}$ of the new parameter sets with the original one, denoted by $\chi_{0}^{2}$. A new parameter set is accepted if it satisfies the criterion,
$\frac{\chi^{2}-\chi_{0}^{2}}{\chi^{2}}<0.2,$ (S5)
i.e., the quality of the description of the ground-state meson spectra remains comparable to that of the original parameter set. Applying this selection criterion, we retain 50 sets of parameters out of around 1500 generated sets. The retained parameter sets are taken to be the variation range of the parameters and are used to estimate the uncertainties. The distribution of the parameters and the meson spectra are shown in Figs. S2 and S3.
Table SM-III: The mean values and standard deviations of multiquark ground-state energies $E$ and binding energies $\Delta E$ calculated using the 50 selected sets of AL1 parameters. For $I(S^{P})=0(1^{+})$ doubly heavy tetraquark systems $T_{cc}$ and $T_{bb}$, we compare the results from the Gaussian Expansion Method (GEM) and DeepQuark (DQ). For $I(S^{P})=0(\frac{5}{2}^{-})$ triply heavy pentaquark systems $P_{cc\bar{c}}$ and $P_{bb\bar{b}}$, the DeepQuark framework is used to solve the ground states.
Systems Method $E$ [MeV] $\Delta E$ [MeV]
$T_{cc}$ GEM $3863\pm 10$ $14.5\pm 0.3$
DQ $3862\pm 10$ $14.9\pm 0.3$
$T_{bb}$ GEM $10477\pm 24$ $150.9\pm 1.0$
DQ $10475\pm 24$ $152.4\pm 1.0$
$P_{cc\bar{c}}$ DQ $5713\pm 14$ $3.1\pm 0.2$
$P_{bb\bar{b}}$ DQ $15544\pm 37$ $13.9\pm 0.3$
We further calculate the multiquark bound states using the 50 selected parameter sets. We calculate the isoscalar vector doubly heavy tetraquark systems $T_{cc}$ and $T_{bb}$ using the GEM and the DeepQuark framework, whose results are shown in Fig. S4. We also calculate ground-state energies of the triply heavy pentaquark bound states $P_{cc\bar{c}}$ and $P_{bb\bar{b}}$ using the DeepQuark framework, whose results are shown in Fig. S5. In Table SM-III we summarize the obtained results, which demonstrate the robustness of our multiquark bound state results and the DeepQuark framework.
  • Within the variation range of the model parameters, the energies of multiquark bound states may vary up to tens of MeV. However, the binding energies $\Delta E$ remain rather stable. The standard deviations of binding energies are around 1 MeV for deeply bound $T_{bb}$ state, and only around 0.3 MeV for shallow bound states $T_{cc},P_{cc\bar{c}},P_{bb\bar{b}}$. This shows the the strength of interactions and the existence of bound states are stable against the changes of parameters.
  • Comparing the DeepQuark results with the GEM results in the doubly heavy tetraquark systems (the first and second row in Fig. S4), we see that the DeepQuark results are in agreement with the GEM results for all 50 parameter sets. This shows that our neural-network framework can consistently describe the ground states of multiquark Hamiltonians with different parameter sets. Moreover, the DeepQuark energy result is slightly lower than the GEM energy result under the same parameter set. The mean values of binding energies is 0.4 MeV lower in the $T_{cc}$ system and 1.5 MeV lower in $T_{bb}$ system, as shown in Table SM-III. This is consistent with the our finding in the manuscript, demonstrating the strong expressive power of DeepQuark.
Figure S2: Distributions of the AL1 model parameter values in the 50 selected parameter sets. The red and green dashed lines represent the mean values over the 50 parameter sets and the original values in Ref. [51], respectively.
Figure S3: Distributions of meson mass spectra (in MeV) calculated using the 50 selected sets of AL1 parameters. The red dashed lines represent the mean values. The red shaded area corresponds to $\pm 1$ standard deviation.
Figure S4: Distributions of ground-state energies $E$ and binding energies $\Delta E$ (in MeV) of the $I(S^{P})=0(1^{+})$ doubly heavy tetraquark bound states, calculated using the 50 selected sets of AL1 parameters. The first and second rows show the results from the DeepQuark (DQ) and Gaussian Expansion Method (GEM), respectively. The red dashed lines represent the mean values. The red shaded area corresponds to $\pm 1$ standard deviation.
Figure S5: Distributions of ground-state energies $E$ and binding energies $\Delta E$ (in MeV) of the $I(S^{P})=0(\frac{5}{2}^{-})$ triply heavy pentaquark bound states, calculated using the 50 selected sets of AL1 parameters. The red dashed lines represent the mean values. The red shaded area corresponds to $\pm 1$ standard deviation.
A.3 Applications to other models
Table SM-IV: Different ingredients in five different quark models.
One-gluon-exchange Confinement meson-exchange
asymptotic freedom Hyperfine form screening effect
AL1 No Gaussian linear No No
AL2 Yes Gaussian linear No No
AP1 No Gaussian $p=2/3$ No No
AP2 Yes Gaussian $p=2/3$ No No
SLM No Yukawa linear Yes Yes
In this work, we present the DeepQuark framework and employ the AL1 model to study the multiquark bound states. However, it should be noted that DeepQuark is not restricted to a specific model, but can be applied to investigate and discriminate among a variety of potential models with different ingredients. To demonstrate this, we employ DeepQuark to calculate the isoscalar vector doubly bottomed tetraquark $T_{bb}$ in five different quark potential models, including the AL1, AP1, AL2, AP2 models introduced in Ref. [51] and the chiral quark model purposed by the group at Salamanca University in Refs. [78, 81] (denoted as SLM). The different ingredients included in the five models are listed in Table SM-IV.
The first four models feature one-gluon-exchange interaction and confinement interaction, taking the form
$V_{ij}=-\frac{3}{16}\boldsymbol{\lambda}_{i}\cdot\boldsymbol{\lambda}_{j}\left(-\frac{\kappa(1-\exp(-r_{ij}/r_{c}))}{r_{ij}}+\frac{8\pi\kappa^{\prime}(1-\exp(-r_{ij}/r_{c}))}{3m_{i}m_{j}}\frac{e^{-r_{ij}^{2}/r_{0}^{2}}}{\pi^{3/2}r_{0}^{3}}\boldsymbol{s}_{i}\cdot\boldsymbol{s}_{j}-\Lambda+\lambda r_{ij}^{p}\right).$ (S6)
The four models differ by the confinement interaction and coupling coefficients. In the AL1 and AL2 models, the power of the confining term is $p=1$, i.e. linear confinement is adopted. In the AP1 and AP2 models, $p=\frac{2}{3}$-power confining term is taken to give the correct asymptotic Regge trajectories. On the other hand, the AL1/AP1 model differs from the AL2/AP2 model by the coupling coefficients of the one-gluon-exchange interaction. In the AL1 and AP1 model, $r_{c}=0$ is taken, namely the coupling coefficients of the Coulomb and hyperfine terms (the first and second terms in Eq. (S6)) are taken to be constants. In the AL2 and AP2 model, $r_{c}\neq 0$ is introduced as a phenomenological simulation of the asymptotic freedom.
The other category of the quark potential model includes one-gluon-exchange interaction, confinement interaction, and the additional one-boson-exchange (OBE) interactions. This is called the chiral quark model, owing to the inclusion of the pseudoscalar meson exchange interactions arising from the spontaneous breaking of chiral symmetry. We choose the SLM model as a representative, whose interactions read
$V_{ij}=\boldsymbol{\lambda}_{i}\cdot\boldsymbol{\lambda}_{j}\left[\frac{\alpha_{s}}{4}\left(\frac{1}{r_{ij}}-\frac{2}{3m_{i}m_{j}}\frac{e^{-r_{ij}/r_{0}}}{r_{0}^{2}r_{ij}}\boldsymbol{s}_{i}\cdot\boldsymbol{s}_{j}\right)+(-a_{c}(1-e^{-\mu_{c}r_{ij}})+\Delta)+V_{ij}^{\text{OBE}}\right].$ (S7)
The regularization of the hyperfine term is taken to be the Yukawa type instead of the Gaussian type in Eq. (S6). In addition, a confinement interaction with color screening effects is adopted, which becomes a linear interaction at short distance and a constant at long distance. The OBE interactions include the pseudoscalar meson exchange stemming from the spontaneous breaking of the chiral symmetry, and the scalar meson exchange, which is used to mimic the two-pion-exchange interaction.
$V_{ij}^{\text{OBE}}=V_{ij}^{\pi}\sum_{a=1}^{3}\left(\lambda_{i}^{f,a}\cdot\lambda_{j}^{f,a}\right)+V_{ij}^{K}\sum_{a=4}^{7}\left(\lambda_{i}^{f,a}\cdot\lambda_{j}^{f,a}\right)+V_{ij}^{\eta}\left[\cos\theta_{P}\left(\lambda_{i}^{f,8}\cdot\lambda_{j}^{f,8}\right)-\sin\theta_{P}\right]+V_{ij}^{\sigma},$ (S8)
with
$V_{ij}^{\chi}=$ $\frac{g_{\mathrm{ch}}^{2}}{4\pi}\frac{m_{\chi}^{2}}{12m_{i}m_{j}}\frac{\Lambda_{\chi}^{2}}{\Lambda_{\chi}^{2}-m_{\chi}^{2}}m_{\chi}\left(\boldsymbol{\sigma}_{i}\cdot\boldsymbol{\sigma}_{j}\right)\left[Y\left(m_{\chi}r_{ij}\right)-\frac{\Lambda_{\chi}^{3}}{m_{\chi}^{3}}Y\left(\Lambda_{\chi}r_{ij}\right)\right],\quad\chi=\pi,K,\eta$ (S9)
$V_{ij}^{\sigma}=$ $-\frac{g_{\mathrm{ch}}^{2}}{4\pi}\frac{\Lambda_{\sigma}^{2}}{\Lambda_{\sigma}^{2}-m_{\sigma}^{2}}m_{\sigma}\left[Y\left(m_{\sigma}r_{ij}\right)-\frac{\Lambda_{\sigma}}{m_{\sigma}}Y\left(\Lambda_{\sigma}r_{ij}\right)\right],$
where $\lambda^{f,a}$ is the Gell-Mann matrix in SU(3) flavor symmetry and $Y(x)=e^{-x}/x$ is the Yukawa function. The model parameters are taken from Ref. [81].
Table SM-V: Ground-state energies (in MeV) of the $I(S^{P})=0(1^{+})$ doubly bottomed tetraquark systems in various quark potential models. The binding energies $\Delta E$ are with respect to the lowest dimeson thresholds $(\bar{B}^{*}\bar{B})$. The ground-state energies obtained from GEM [60] are listed for comparison.
Potential DQ GEM    [60]
$E$ $\Delta E$ $E$ $\Delta E$
AL1 10491 -153 10493 -152
AP1 10502 -176 10504 -175
AL2 10473 -151 10474 -151
AP2 10471 -174 10472 -174
SLM 10230 -362 10232 -360
We calcualte the ground-state energies of $T_{bb}$ in the five models, and compare the DeepQuark results with the GEM results in Table SM-V. We find that DeepQuark can be successfully applied to all the models considered above, yielding ground-state energies that are consistent with and slightly lower than the corresponding GEM results. This demonstrates that DeepQuark can accommodate a wide variety of interactions, including different confining mechanisms and short-range dynamics, without being tied to a specific model. Combined with its capability to handle complex multiquark systems such as pentaquark states, DeepQuark provides a general and flexible framework for phenomenological studies of multiquark spectrum.
Supplemental Material B Details of the optimization process